Project

General

Profile

Actions

Agenda

  • Actions
    • Macwaters - Remove old overheated CSW71 to R9 waiting for John Govans and the van - in progress
    • Hall Manager - water system change-over procedure (stands) - collar people on Friday
    • Progress on Hall list MICE-Work-List-May-2017.pdf (Colin) Remaining items_
    • AKurup to check which settings are being used by the active Alarm Handlers
    • CWhyte to request survey of CC Magnets after cycle
    • FDrielsma to investigate low-efficiency channels in EMR x-z
    • Tracker group: decide what should be done to prevent repeat contamination of Tracker Cryostats
  • DC update
  • Hydrogen system update and plans
  • Issues to fix before September : Cycle 2017 Issues
  • Mobile Logistics and Operational Support update (CWhyte)
  • AOB

Minutes

  • Actions
    • Macwaters - Remove old overheated CSW71 to R9 waiting for John Govans and the van - in progress
    • Hall Manager - water system change-over procedure (stands) - have prescription, need document (Colin?)
    • Progress on Hall list MICE-Work-List-May-2017.pdf (Colin) Remaining items_
      • SS: stands. FC/LH2: notes below. Crate fan checks: need care as we've frequently had crates fail before, but note that we saw alarms from a number of crates during the last run.
    • AKurup to check which settings are being used by the active Alarm Handlers
    • CWhyte to request survey of CC Magnets after cycle - stands
    • FDrielsma to investigate low-efficiency channels in EMR x-z investigated, nothing he can do
    • Tracker group: decide what should be done to prevent repeat contamination of Tracker Cryostats - Wednesday mtg.
    • Status of miceraid5 disk cleanup (stands) EOverton now has a script but needs to run it: needs to hand over the procedure and script to someone, i.e. Durga
  • DC update
    Handover from Operations last Thursday. The LH2 control system is unstable when there is no liquid at all present. The FC quench test with LH2 Absorber is to happen tomorrow.
    The Decay Solenoid has been warmed up for its service (2 days) but Joerg is busy elsewhere at RAL today; will try tomorrow. Mark Tucker is the liaison.
    Conventional magnet water flow has been turned down to allow conductivity to rise.
    During shutdown will need to look at the Target. Will also open up the DSA.
    LH2 - see below.
    No news regarding John Govans.
  • Hydrogen system update and plans
    The FC will be ramped and quenched tomorrow: need to check the Quench Detection logging. John Cobb will come tomorrow.
    The LH2 system presently contains some liquid Neon - controls are unstable with gas only. Have lost 30 bar of Neon from the bottle somewhere.
    The second LH2 turret has been brazed; needs leak-checking and to have the shield fitted. Shield still needs mods.
    LH2 controls data will be converted to allow archiving within the EPICS Archiver.
  • Issues to fix before September : Cycle 2017/01 Issues
    Heated debate about whether the C&M system was in a fit state for a production data-taking run.
    HNebrensky and AKurup will work through the C&M items on the list. Others can be assigned to experts but who will be responsible overall remains unclear.
  • Mobile Logistics and Operational Support update (CWhyte)
    • Yes.
  • AOB
    The Trackers still have a lot of dead channels, so will need to move waveguide connections around in the next couple of weeks - Ed will need access to the Hall and help from Geoff Barber and Craig.

Updated by Nebrensky, Henry almost 6 years ago · 12 revisions